# Studying stellar populations in Omega Centauri with phylogenetics

### v1

@@ 3.1 Stellar Population Barplot
Given that we have found that our data is chemically diverse, we proceed to cluster the stars using all abundance ratios, [X/Fe], with Gaussian Mixture Models (GMM). The Bayesian Information Criterion (BIC) scores indicate that the optimal number of units or groups in this dataset is three (see Appendix. B.2). In Fig. 5 we show the GMM classification for 3 groups for all stars. This plot is similar to the STRUCTURE barplot, which is a popular tool for studying population structure in biology (Pritchard et al., 2000). While STRUCTURE uses a mixture model for DNA, which is a categorical variable, we adapt the idea to describe stellar populations by using GMM for chemical abundance ratios. We call our approach Stellar Population Barplot.
@@
Figure 5: Stellar Population Barplot. Each star, whose name is indicated in the X-axis, is shown as a bar with the cumulative probability of the star belonging to a specific group. The Y-axis shows the probability of the star belonging to the groups, which are assigned using a Gaussian Mixture Model of 3 components. The Stellar Population Barplot allows us to appreciate the clustering, and to identify which groups might be more related to each other, as reflected by the level of mixing between them.
@@
The main difference between the Stellar Population Barplot and how astrophysicists usually use GMM (Das et al., 2020; Buckley et al., 2024; Buder et al., 2022) is that we do not only select stars with high probability of belonging to a certain GMM group for further analyses. Instead, the barplot includes stars with all levels of probabilities. This is because stars with low probability of belonging to a certain group offers us the opportunity to study mixing between groups (Villarreal et al., 2022; Nespolo et al., 2020). This is similar to the STRUCTURE barplot where low probability can be indicative of migration or admixture events, e.g. when a particular population is descended from multiple ancestral populations (Peter, 2016). If the majority of stars have a 100% probability of belonging to one GMM group, then the distribution of chemical abundances over stars is discrete, suggesting that each GMM group is likely to have evolved independently. But if the majority of stars have a low probability of belonging to a single group, then the distribution of chemical abundances over stars is continuous. This suggests strong mixing among groups, for example, the chemical elements from one group is mixed with the interstellar medium that forms the other group.
@@
In our case, the orange Group_O1 has only a handful of stars from the purple Group_O3. This would mean that the orange group has evolved independently to the other groups. On the other hand, the gray group has a handful of stars with more than 20% of contribution from the orange and purple groups, suggesting that the gray group could be more related to the others. This interpretation is in any case subject to the uncertainties of the abundance measurements, as well as the latent chemical space considered for the clustering (Buder et al., 2022; Buckley et al., 2024), and also the number of GMM groups. To draw firm conclusions, we would need validate with simulated data, such as Gray et al. (2024), where the history is known. Various key abundance ratio planes color-coded by the GMM groups can be found in Appendix. B.3, where they are discussed at length.
@@ 3.2 Phylogenetic tree of $\omega$ Cen
Figure 6: Circular Optical tree. The colors of the tip labels correspond to the GMM groups. The pie charts on the nodes correspond to the percentage of occurrence of the node in 1000 random generated trees considering the abundance uncertainties, indicating that the tree has an overall support of 35%.
@@ 3.2.1 Building NJ trees
Following de Brito Silva et al. (2024), we built trees using the empirical distance method and the agglomerative neighbor-joining (NJ) clustering algorithm (Saitou & Nei, 1987), which does not rely on an evolutionary model to find the best tree from a given dataset.
This is important when borrowing tools from a different discipline, as we must avoid making assumptions that might not apply to astrophysics. The NJ algorithm arranges chemically similar stars nearer to one another in the tree while placing those more different further apart. The branch lengths correspond to the overall differences between stars. This information is taken from a distance matrix, which is symmetric and of the dimension of the number of stars. In this case, we computed Manhattan distances based on stellar abundances to determine these differences. The pairwise distance matrix was calculated using the [X/H] abundances ratios available to us, namely [O/H], [Na/H], [Al/H], [Si/H], [Ca/H], [Sc/H], [Ti/H], [Fe/H], [Ni/H], [La/H] and [Eu/H].
To assess the robustness of our tree topology we calculate the node support by considering the uncertainties in our abundances. We sampled 1000 distances matrices, by perturbing the abundances with their uncertainties. From these distance matrices we calculated 1000 NJ trees. We counted how many of these trees contained each node of the original one.
@@ 3.2.2 Stellar phylogenies
Figure 6 shows the tree in circular form and allows us to realize that the tree shows similar results to GMM in the sense that purple and grey stars form two ‘caterpillar’ trees, suggesting they evolved independently. They are highlighted in the yellow and cyan boxes. The remaining stars, highlighed in the pink box, form a more ‘star-like’ tree, suggesting that they either mixed a lot or are from different sources. The nodes of the figure have pie-charts overplotted, which indicate the the proportion of trees that have this node in blue given the uncertainties in the abundance measurements. Due to uncertainties in abundance measurements, multiple pair-wise distances between stars in their chemical composition are possible, causing tree topologies to differ among each other. The overall support of 35% is similar to Walsen et al. (2024), which is expected since this is the support due to uncertainties in the stellar chemical data.
When the root, e.g. the ancestral node, is set, and assuming the choice of root is sensible (Yang, 2014), the phylogenetic tree reveals more information on star formation history. The resulting rooted tree is displayed in Fig. 7. We set the root at the star with the lowest [Ca/H] abundance ratio. This assumes that chemical abundance increases with time, so the lowest abundances in a dataset should reflect the earliest time and therefore a root for the tree. We choose [Ca/H] instead of [Fe/H] in this case because of its excellent consistency in the abundance measurements with the Infrared sample, which allows us to use the same criterion to root the Infrared tree. Furthermore, the PCA revealed that Ca had a similar effect on the chemical diversity of both samples. This enables a better comparison between trees and so a better interpretation of our results (see Sect. 4.2). Our analysis revealed that stars with pink labels stars, i.e. stars whose probability to belong to a specific group is lower than 90%, do not have a preferred location in the tree. Changing the threshold for this definition does not affect our interpretation.
@@ 4 The assembly of populations in $\omega$ Cen
In this section we interpret the astrophysical meaning of the branches and groups found in our phylogenetic analysis, and so discuss the possible history of assembly of $\omega$ Cen.
@@ 4.1 The Block, Child and GC branches of the tree
Figure 6 and 7 show two big branches, each with a clear ‘caterpillar’ topology. We call these two branches Block and Child. The remaining tips are closer to the root of the tree and show a more star-like topology, which we call GC. Considering our null hypothesis (de Brito Silva et al., 2024), the Block and Child branches likely evolved independently, while group GC likely consists of stars well mixed or from different sources. The chemical patterns in Block, Child, and GC groups are not only distinct, but also suggest their possible history. The next four panels to the right of Fig. 7 show the distribution of [Fe/H], [O/Fe], [Al/Fe] and [La/Fe] along the branches. Evidence supporting Block and Child having independent star formation histories is that the two branches have different [O/Fe], [Al/Fe] and [La/Fe] trends. Also, the majority of purple stars, grouped by GMM, are in the Block branch and the majority of gray stars are in the Child branch. In contrast, GC stars have a narrow [Fe/H] and [La/Fe] ranges, and present the Al-O, Na-O anticorrelation which is typical in globular clusters, and so we call this population GC-Cen.
@@
The population of stars from the Block branch shows a classical [O/Fe] trend with metallicity, namely the [O/Fe]-[Fe/H] relation presenting a potential [O/Fe] knee at $\mathrm{[Fe/H]}\sim-1.3$. This is consistent with the chemical evolution of a dwarf galaxy (Matteucci & Brocato, 1990). The increasing trend of [La/Fe] with [Fe/H] is also found in dwarf galaxies. Since La is a neutron-capture element produced via the slow-process mechanism in AGB stars, reaching high [La/Fe] abundance ratios is possible only after a few Gyr (Bisterzo et al., 2014; Romano et al., 2023). [Al/Fe] is notably more enhanced than in typical dwarf galaxies, which are normally subsolar and lower compared to the Milky Way trend (Das et al., 2020; Hasselquist et al., 2021). We note however that recent discoveries of Milky Way metal-poor ancient structures, such as Shiva and Shakti (Malhan & Rix, 2024), as well as the G3/G5 groups from the GALAH sample by Buckley et al. (2024), have significantly higher [Al/Fe] than the better-known disrupted galaxies in the halo of comparable metallicities. Such systems were proposed by Malhan & Rix (2024) to have formed from gas clumps in a massive and/or dense progenitor. The abundance patterns of the Block branch could be thus associated with a building block population a proto-galaxy, equivalent to Shiva and Shakti of the Milky Way. We thus call this population Block-Cen and attribute it to the building block of the progenitor galaxy that formed $\omega$ Cen.
@@
The Child branch has an anti-correlation between Al and O like the GC branch, although [Fe/H] has a significant increasing trend, as does [La/Fe]. This point towards an extended star formation history, perhaps a result of an in-situ star formation from material that has been enriched by the stars within $\omega$ Cen, henceforth its designation as Child-Cen. Indeed, the G_O2-gray stars has slightly more admixture with G_O1-orange and G_O3-purple in the Stellar Population Barplot, suggesting that this group is somehow more related to the other populations. This admixture is driven by [Ca/Fe] and [La/Fe] abundances (see Fig. 15). The abundance patterns of Child-Cen, namely the very low [O/Fe] and high [Al/Fe], have not been seen in the local Universe. This suggests that the Child stars could be the product of a different star formation environment to Block-Cen or any other population found in the Milky Way.
Figure 7: Neighbor-Joining tree of $\omega$ Cen stars. Tips are coloured based on the GMM group to which each star has the highest probability of belonging to in Fig. 5. The tree shows 3 main branches, whose [Fe/H], [O/Fe], [Al/Fe], and [La/Fe] are shown in the right-hand panels. The GC branch has properties similar to globular clusters. The Block and Child branches have extended metallicity and [La/Fe] distributions, supporting in-situ star formation for a prolonged duration. However, their [O/Fe] and [Al/Fe] trends are distinct, suggesting that they had different star formation histories and explaining why they are placed on different branches and GMM groups.
@@ 4.2 Assessing systematic differences with common stars
To understand further the similarities and the differences in the populations and branches found between both datasets (see Appendix. C for the tree obtained from the Infrared sample), as well as the tree shapes, we build trees of the stars in common between both datasets. The trees are seen in Fig. 8. The left hand trees correspond to the Optical dataset and the right hand ones to the Infrared dataset. Same stars are connected through lines. The top trees are a subset of the trees from Fig. 7 and Fig. 17, i.e. they are built using all abundance ratios but only for the 16 stars of Tab. 1. Rooting the tree by the star with lowest [Ca/H], which is one of the elements with best agreement between both datasets, allowing us to use the same root for both trees. But the topology does not agree among both datasets, as seen by the various crossed lines. These indicate the conflict of branches, which gives us a warning of the dependency of the elements used in order to reconstruct the history of $\omega$ Cen using these trees. This conflict of branches thus causes the GC/Block blue and the GC red branches of the Infrared tree to be placed at locations that are not expected from the Optical tree.
@@
To quantify the difference, we calculate the normalized Robinson-Foulds distances ($\mathrm{RFD}$), following de Brito Silva et al. (2024). $\mathrm{RFD}=0$ means trees have identical branching pattern, and $\mathrm{RFD}=1$ means trees have totally different branching patterns. An extensive explanation of this distance measure and its pros and cons in using it can be found in Yang (2014). Between the two upper trees of Fig. 8, we obtain $\mathrm{RFD}=0.63$. The lower panels of Fig. 8 are trees built using only the common abundance between the two (see the elements plotted in Fig. 3). The ranking of the tips between both trees agrees better, but the trees still differ by $\mathrm{RFD}=0.57$. There are only a few stars that are strongly disordered, with 38% of the stars ordered identically. And, if we remove SOP_29/SIR_58, that increases to 60%. This seems like a large improvement on the all-abundance tree comparison, but the RFD score remains quite similar. This might be due to the uncertainties of the abundance measurements which are driving significant scatter in the tree shapes, as noted by the averaged 35% of node support discussed above. However, we need to consider the limitations of the RFD distance metric, which only focuses on the topology of the tree and counts the number of common nodal position between trees (see Yang, 2014). We discuss this issue further below.
@@
It is actually interesting that these $\mathrm{RFD}$ are not significantly different when using the full abundance dataset or only the common abundances. In order to assess if this similarity is significant given the uncertainties of the measurements, we can estimate the minimum expected difference for the same tree, given its uncertainties. To do so, we run another Monte Carlo simulation to calculate random distance matrices considering the abundances and the reported uncertainty. We then build NJ trees and compare them among each other. We do this 1000 times and estimate the mean $\mathrm{RFD}$ and its standard deviation for the Optical and for the Infrared sample independently. This was done using the common abundances between two samples, as well as the full set of abundances.
@@
For the full set of abundances, that is Fe, O, Na, Al, Si, Ca, Sc, Ti, Ni, La and Eu for the Optical and Fe, O, Mg, Al, Si, Ca and Ce for the Infrared, we obtain that trees are similar among each other on average by $\mathrm{RFD}=0.31\pm 0.11$ and $\mathrm{RFD}=0.35\pm 0.1$, respectively. These numbers are similar to each other, and comparable to the 35% of the nodal support. In both cases, adding abundances increases the robustness of the tree topology accounting for abundance measurement errors, since $\mathrm{RFD}$ decreases. We note that the improvement of the Infrared $\mathrm{RFD}$ is very small. This is because only two additional elements, namely Mg and Ce, have been added to the distance matrix. The value of $\mathrm{RFD}=0.63$ for the difference of the Optical and the Infrared tree using all the abundances displayed in the upper panel of Fig. 8 is also the result of more precise trees and therefore the difference can not be explained only due to errors.
It is worth referring to to the discussion of Kuhner & Yamato (2014). The metric of RFD counts the number of branch partitions that appear in one tree but not the other, scoring 1 for each non-matched partition. This means that the $\mathrm{RFD}$ metric does not consider differences in branch lengths, only tree topology. Furthermore, a rooted-tree has a maximum of number of nodes of $2(n-2)$, where $n$ is the number of tips. This means $\mathrm{RFD}$ has $2(n-2)$ partitions to score. In our case, $n=16$ implies we score 28 partitions only, which is not a very large number to quantify significant differences between the trees here.
Even considering these caveats, we obtain that Optical and Infrared trees are different to each other. This difference is dominated by the systematic uncertainties of the measurements rather than the election of elements used in the distance matrix. A reason for this result could be that even if the full Optical and Infrared sample consider different elements, these elements represent the same 3 nucleosynthetic channels, namely, CCSNe, SNe-Ia, and AGB winds. In both datasets $\alpha-$capture, iron-peak and neutron-capture elements are included. Indeed, by removing e.g. Fe from the Infrared sample, the $\mathrm{RFD}$ between the Optical and the Infrared sample increases to a mean of $\mathrm{RFD}=0.92\pm 0.01$. Removing Ce has a much smaller impact on the $\mathrm{RFD}$ estimate. This could be related to the fact that Fe is a more significant driver to the variance in the abundance space compared to Ce, as discussed in Sect. B.1.

### v2

@@ 3.1 Stellar population bar plot
Given that we have found that our data is chemically diverse, we proceed to cluster the stars using all abundance ratios, [X/Fe], with Gaussian mixture models (GMM). The Bayesian Information Criterion (BIC) scores indicate that the optimal number of units or groups in this dataset is three (see Appendix. B.2). In Fig. 4 we show the GMM classification for three groups for all stars. This plot is similar to the STRUCTURE bar plot, which is a popular tool for studying population structure in biology (Pritchard et al., 2000). While STRUCTURE uses a mixture model for DNA, which is a categorical variable, we adapt the idea to describe stellar populations by using GMM for chemical abundance ratios. We call our approach stellar population bar plot.
@@
Figure 4: Stellar population bar plot. Each star, whose name is indicated in the X axis, is shown as a bar with the cumulative probability of the star belonging to a specific group. The Y axis shows the probability of the star belonging to the groups, which are assigned using a Gaussian mixture model of three components. The stellar population bar plot allows us to appreciate the clustering, and to identify which groups might be more related to each other, as reflected by the level of mixing between them.
@@
The main difference between the stellar population bar plot and how astrophysicists usually use GMM (Das et al., 2020; Buckley et al., 2024; Buder et al., 2022) is that we do not only select stars with high probability of belonging to a certain GMM group for further analyses. Instead, the bar plot includes stars with all levels of probabilities. This is because stars with a low probability of belonging to a certain group offers us the opportunity to study mixing between groups (Villarreal et al., 2022; Nespolo et al., 2020). This is similar to the STRUCTURE bar plot where low probability can be indicative of migration or admixture events, for example when a particular population is descended from multiple ancestral populations (Peter, 2016). If the majority of stars have a 100% probability of belonging to one GMM group, then the distribution of chemical abundances over stars is discrete, suggesting that each GMM group is likely to have evolved independently. But if the majority of stars have a low probability of belonging to a single group, then the distribution of chemical abundances over stars is continuous. This suggests strong mixing among groups, for example, the chemical elements from one group is mixed with the interstellar medium that forms the other group.
@@
Most of the stars with a high probability of being in Group_O1 (orange) have very low probabilities of being in another group. This suggests that Group_O1 has evolved independently of the other groups. This independency might be spatially or temporally, namely this group could have formed far away from the other groups and/or before the other groups existed. On the other hand, the gray group has a handful of stars with more than 20% of contribution from the orange and purple groups, suggesting that the gray group could be more related to the others. This interpretation is in any case subject to the uncertainties of the abundance measurements, as well as the latent chemical space considered for the clustering (Buder et al., 2022; Buckley et al., 2024), and also the number of GMM groups. To draw firm conclusions, we would need validate with simulated data, such as Gray et al. (2025), where the history is known. Various key abundance ratio planes color-coded by the GMM groups can be found in Appendix. B.3, where they are discussed at length.
@@ 3.2 Phylogenetic tree of $\omega$ Cen
Figure 5: Circular Optical tree. The colors of the tips correspond to the GMM groups. The pie charts on the nodes correspond to the percentage of occurrence of the node in 1000 random generated trees considering the abundance uncertainties (node support), indicating that the tree has an overall support of 35%.
@@ 3.2.1 Building NJ trees
Following de Brito Silva et al. (2024), we built trees using the empirical distance method and the agglomerative neighbor-joining (NJ) clustering algorithm (Saitou & Nei, 1987), which does not rely on an evolutionary model to find the best tree from a given dataset. This is important when borrowing tools from a different discipline, as we must avoid making assumptions that might not apply to astrophysics. The NJ algorithm arranges chemically similar stars nearer to one another in the tree while placing those more different further apart. The branch lengths correspond to the overall chemical differences between stars (see also Sect. 3 of Jofré et al., 2017). The chemical distance is taken from a distance matrix, which is symmetric and of the dimension of the number of stars. In this case, following Jofré et al. (2017), we computed Manhattan distances based on stellar abundances. The pairwise distance matrix was calculated using the [X/H] abundance ratios available to us, namely [O/H], [Na/H], [Al/H], [Si/H], [Ca/H], [Sc/H], [Ti/H], [Fe/H], [Ni/H], [La/H], and [Eu/H].
To assess the robustness of our tree topology we calculate the node support by considering the uncertainties in our abundances. We sampled 1000 distance matrices, by perturbing the abundances with their uncertainties. From these distance matrices we calculated 1000 NJ trees. The node support corresponds to the percentage of occurrence of a given node in all other sampled trees. This is calculated by counting the occurrence of nodes in all trees (see Jackson et al., 2021; Walsen et al., 2024, for more details about node support in stellar trees).
@@ 3.2.2 Stellar phylogenies
Figure 5 shows the tree in circular form and allows us to realize that the tree shows similar results to the GMM analysis in the sense that purple and gray stars form two ‘caterpillar’ trees, suggesting they evolved independently. Since these branches contain stars from the different GMM groups, we have labeled with new names, which are explained in detail in the next Section. The branches are highlighted in the yellow and cyan boxes. The remaining stars, highlighed in the pink box, form a more ‘star-like’ tree, suggesting that they either mixed a lot, are from different sources, or the data is too uncertain for the NJ method to resolve the hierarchical differences there. The nodes of the figure have pie charts overplotted, which indicate the proportion of trees that have this node in blue given the uncertainties in the abundance measurements (this is the node support). Due to uncertainties in abundance measurements, multiple pair-wise distances between stars in their chemical composition are possible, which translates in an uncertain tree topology. It is possible to see that the orange stars, which belong to the Group_O1 from the GMM and are mostly part of the pink and the deep nodes of the Block box, have nodes with low support. This is because these stars are so similar to each other that the NJ algorithm is not able to assign them to a specific node given the uncertainties. Indeed, here we find ourselves with a limitation of the NJ method, already discussed in Walsen et al. (2024). By construction, the NJ has to assign every star into a different branch, assuming each of them represents a different star formation episode. When analyzing observed stars, we cannot rule out the possibility that two stars belong to the same star formation episode. While the GMM will put them together in the same group (because they might have the same chemical signature), the NJ algorithm has to place them in distinct leaves. The overall support of 35% is similar to Walsen et al. (2024), which is expected since this is the support due to uncertainties in the stellar chemical data.
@@ 4 The assembly of populations in $\omega$ Cen
When the root, that is, the ancestral node, is set, and assuming the choice of root is sensible (Yang, 2014), the phylogenetic tree reveals more information on star formation history. The resulting rooted tree is displayed in Fig. 6. We set the root at the star with the lowest [Ca/H] abundance ratio. This assumes that chemical abundances increases with time, so the lowest abundances in a dataset should reflect the earliest time and therefore a root for the tree. We choose [Ca/H] instead of [Fe/H] in this case because of its excellent consistency in the abundance measurements with the Infrared sample, which allows us to use the same criterion to root the Infrared tree. Furthermore, the PCA revealed that Ca had a similar effect on the chemical diversity of both samples. This enables a better comparison between trees and so a better interpretation of our results (see Sect. 4.2).
Figure 6: Neighbor-joining tree of $\omega$ Cen stars. Tips are colored based on the GMM group to which each star has the highest probability of belonging to in Fig. 4. The tree shows three main branches, whose [Fe/H], [O/Fe], [Al/Fe], and [La/Fe] are shown in the right-hand panels. The GC branch has properties similar to globular clusters. The Block and Child branches have extended metallicity and [La/Fe] distributions, supporting in-situ star formation for a prolonged duration. However, their [O/Fe] and [Al/Fe] trends are distinct, suggesting that they had different star formation histories and explaining why they are placed on different branches and GMM groups.
Our analysis revealed that stars with pink labels stars, that is, stars whose probability to belong to a specific group is lower than 90%, do not have a preferred location in the tree. Changing the threshold for this definition does not affect our interpretation. In this section we interpret the astrophysical meaning of the branches and groups found in our phylogenetic analysis, and so discuss the possible history of assembly of $\omega$ Cen.
@@ 4.1 The Block, Child and GC branches of the tree
. Figure 5 and 6 show two big branches, each with a clear ‘caterpillar’ topology. We call these two branches Block and Child. The remaining tips are closer to the root of the tree and show a more star-like topology, which we call GC. Considering our null hypothesis (de Brito Silva et al., 2024), the Block and Child branches likely evolved independently, while group GC likely consists of stars well mixed or from different origins. The chemical patterns in Block, Child, and GC groups are not only distinct, but also suggest their possible history. The next four panels to the right of Fig. 6 show the distribution of [Fe/H], [O/Fe], [Al/Fe], and [La/Fe] along the branches. Evidence supporting Block and Child having independent star formation histories is that the two branches have different [O/Fe], [Al/Fe], and [La/Fe] trends. Also, the majority of purple stars, grouped by GMM, are in the Block branch and the majority of gray stars are in the Child branch. In contrast, GC stars have a narrow [Fe/H] and [La/Fe] ranges, and present the Al-O, Na-O anticorrelation, which is typical in globular clusters, and so we call this population GC-Cen.
@@
The population of stars from the Block branch shows a classical [O/Fe] trend with metallicity, namely the [O/Fe]-[Fe/H] relation with a [O/Fe] plateau at 0.4 dex, which is the signature of SNII and presenting a potential [O/Fe] knee at $\mathrm{[Fe/H]}\sim-1.3$, which results from the pollution of [Fe/H] due to SNIa. This is consistent with the chemical evolution of a dwarf galaxy (Matteucci & Brocato, 1990). The increasing trend of [La/Fe] with [Fe/H] is also found in dwarf galaxies. Since La is a neutron-capture element produced via the slow-process mechanism in AGB stars, reaching the high [La/Fe] abundance ratios is possible only after a few Gyr (Bisterzo et al., 2014; Romano et al., 2023). [Al/Fe] is notably more enhanced than in typical dwarf galaxies, which are normally subsolar and lower compared to the Milky Way trend (Das et al., 2020; Hasselquist et al., 2021). We note however that recent discoveries of Milky Way metal-poor ancient structures, such as Shiva and Shakti (Malhan & Rix, 2024), as well as the G3/G5 groups from the GALAH sample by Buckley et al. (2024), have significantly higher [Al/Fe] than the better-known disrupted galaxies in the halo of comparable metallicities. Such systems were proposed by Malhan & Rix (2024) to have formed from gas clumps in a massive and/or dense progenitor. The abundance patterns of the Block branch could be thus associated with a building block population or a proto-galaxy, equivalent to Shiva and Shakti of the Milky Way. We thus call this population Block-Cen and attribute it to the building block of the progenitor galaxy that formed $\omega$ Cen.
@@
The Child branch has an anticorrelation between Al and O like the GC branch, although [Fe/H] has a significant increasing trend, as does [La/Fe]. This points towards an extended star formation history, perhaps a result of in-situ star formation from material that has been enriched by the stars within $\omega$ Cen, henceforth its designation as Child-Cen. Indeed, the G_O2-gray stars has slightly more admixture with G_O1-orange and G_O3-purple in the stellar population bar plot, suggesting that this group is somehow more related to the other populations. This admixture is driven by [Ca/Fe] and [La/Fe] abundances (see Fig. 14).
The anticorrelations of stellar abundances such as [O/Fe] and [Al/Fe] are among the most challenging features of globular clusters to be explained (see Gratton et al., 2019, for extensive review). While there are strong arguments in favor of an inheritance process due to the CNO and MgAl cycles inside massive stars, which pollute the ISM in the AGB phase via winds (D’Ercole et al., 2012; Gratton et al., 2019), there are still many theoretical aspects on stellar nucleosynthesis and GC formation that are not understood and therefore the observations of such abundance ratios cannot be fully reproduced for all GCs (Bastian & Lardo, 2018). The case of $\omega$ Cen offers additional complexities due to the extended range of [Fe/H] and s-process elements, which suggest timescales of Gyr for star formation compared to the few Myr of the CNO, NeNA, and MgAl cycles of massive stars. The peculiar abundance patterns of Child-Cen, namely the very low [O/Fe], and high [Al/Fe], cannot be directly associated with the (unexplained) O-Al anticorrelation of classical GCs, because of the different timescales needed to simultaneously describe the pollution mechanisms of the anticorrelations and the chemical enrichment of the Fe and La trends. A stellar population with such characteristics has not been found in the Galactic field, or in mono-metallic clusters. The chemical pattern of Child-Cen would therefore need a new explanation.
@@ 4.2 Assessing systematic differences with common stars
To understand further the similarities and the differences in the populations and branches found between both datasets (see Appendix. C for the tree obtained from the Infrared sample), as well as the tree shapes, we build trees of the stars in common between both datasets. The trees are seen in Fig. 7. The left hand trees correspond to the Optical dataset and the right hand ones to the Infrared dataset. Same stars are connected through lines. The top trees are a subset of the trees from Fig. 6 and Fig. 16, that is, they are built using all abundance ratios but only for the 16 stars of Tab. 1. Rooting the tree by the star with lowest [Ca/H], which is one of the elements with best agreement between both datasets, allows us to use the same root for both trees. But the topology does not agree among both datasets, as seen by the various crossed lines. This suggests the choice of abundances ratios used affects the order of the tips in the tree. This conflict of branches thus causes the GC/Block blue and the GC red branches of the Infrared tree to be placed at locations that are not expected from the Optical tree.
@@
To quantify the difference, we calculate the normalized Robinson-Foulds distances ($\mathrm{RFD}$), following de Brito Silva et al. (2024). $\mathrm{RFD}=0$ means trees have identical branching pattern, and $\mathrm{RFD}=1$ means trees have totally different branching patterns. An extensive explanation of this distance measure and its pros and cons in using it can be found in Yang (2014). Between the two upper trees of Fig. 7, we obtain $\mathrm{RFD}=0.63$. The lower panels of Fig. 7 are trees built using only the common abundance between the two (see the elements plotted in Fig. 2). The ranking of the tips between both trees agrees better, but the trees still differ by $\mathrm{RFD}=0.57$. There are only a few stars that are strongly disordered, with 38% of the stars ordered identically. And, if we remove SOP_29/SIR_58, that increases to 60%. This seems like a large improvement on the all-abundance tree comparison, but the RFD score remains quite similar. This might be due to the uncertainties of the abundance measurements, which are driving significant scatter in the tree shapes, as noted by the averaged 35% of node support discussed above. However, we need to consider the limitations of the RFD distance metric, which only focuses on the topology of the tree and counts the number of common nodal position between trees (see Yang, 2014). We discuss this issue further below.
@@
It is actually interesting that these $\mathrm{RFD}$s are not significantly different when using the full abundance dataset or only the common abundances. In order to assess if this similarity is significant given the uncertainties of the measurements, we can estimate the minimum expected difference for the same tree, given its uncertainties. To do so, we run another Monte Carlo simulation to calculate random distance matrices considering the abundances and the reported uncertainty. We then build NJ trees and compare them. We do this 1000 times and estimate the mean $\mathrm{RFD}$ and its standard deviation for the Optical and for the Infrared sample independently. This was done using the common abundances between two samples, as well as the full set of abundances.
@@
For the full set of abundances, that is Fe, O, Na, Al, Si, Ca, Sc, Ti, Ni, La, and Eu for the Optical and Fe, O, Mg, Al, Si, Ca, and Ce for the Infrared, we obtain that trees are similar among each other on average by $\mathrm{RFD}=0.31\pm 0.11$ and $\mathrm{RFD}=0.35\pm 0.1$, respectively. These numbers are similar to each other, and comparable to the 35% of the nodal support. In both cases, adding abundances increases the robustness of the tree topology accounting for abundance measurement errors, since $\mathrm{RFD}$ decreases. We note that the improvement of the Infrared $\mathrm{RFD}$ is very small. This is because only two additional elements, namely Mg and Ce, have been added to the distance matrix. The value of $\mathrm{RFD}=0.63$ for the difference of the Optical and the Infrared tree using all the abundances displayed in the upper panel of Fig. 7 is also the result of more precise trees and therefore the difference can not be explained only due to errors.
It is worth referring to to the discussion of Kuhner & Yamato (2014). The RFD metric counts the number of branch partitions that appear in one tree but not the other, scoring one for each non-matched partition. This means that the $\mathrm{RFD}$ metric does not consider differences in branch lengths, only tree topology. Furthermore, a rooted-tree has a maximum of number of nodes of $2(n-2)$, where $n$ is the number of tips. This means $\mathrm{RFD}$ has $2(n-2)$ partitions to score. In our case, $n=16$ implies we score 28 partitions only, which is not a very large number to quantify significant differences between the trees here.
Even considering these caveats, we find that the Optical and Infrared trees are different to each other. This difference is dominated by the systematic uncertainties of the measurements rather than the election of elements used in the distance matrix. A reason for this result could be that even if the full Optical and Infrared sample consider different elements, these elements overlap in three nucleosynthetic channels, namely, CCSNe, SNe-Ia, and AGB winds. In both datasets $\alpha-$capture, iron-peak, and neutron-capture elements are included. We note that the Optical sample includes Eu, which is an r-process element produced by neutron star mergers, but Eu has a weaker contribution to the variance of the Optical data, according to the PCA (see Sect. B.1). Indeed, by removing for instance Fe from the Infrared sample, the $\mathrm{RFD}$ between the Optical and the Infrared sample increases to a mean of $\mathrm{RFD}=0.92\pm 0.01$. Removing Ce has a much smaller impact on the $\mathrm{RFD}$ estimate. This could be related to the fact that Fe is a more significant driver of the variance in the abundance space compared to Ce, as discussed in Sect. B.1.